OpenStudy (anonymous):

\[What's~~~f'~~~ of ~~~ \sin^2x^2+\cos^2x^2~~~~~~?\]

OpenStudy (anonymous):

In other words, differentiate it.

OpenStudy (anonymous):

I think I got a wrong answer... I got 4sinx^2cosx^2 +4cosx^2sinx^2

OpenStudy (anonymous):

Or, is it -4cosx^2sinx^2?

OpenStudy (mani_jha):

\[2sinx ^{2}\cos x^{2}2x+2cosx ^{2}(-\sin x^{2})2x\]

OpenStudy (mani_jha):

Yes it is -4cosx^2sinx^2

OpenStudy (anonymous):

My text book says the answer is 0? Is it wrong?

sam (.sam.):

actually its 0

sam (.sam.):

wolfram got problem, sin^(2)x^(2)+cos^(2)x^(2) (d)/(dv) =cos(v) (d)/(dx) =4xsin((x^(2)))cos((x^(2))) (d)/(dv) =-sin(v) (d)/(dx) =-4xcos((x^(2)))sin((x^(2))) (d)/(dx) =4xsin((x^(2)))cos((x^(2)))-4xcos((x^(2)))sin((x^(2))) 0

OpenStudy (anonymous):

Oh. So, if you minus what you and I differentiates, the answer is 0?

OpenStudy (anonymous):

differentiated*

sam (.sam.):

yes

OpenStudy (anonymous):

Ok, great :) Thanks